OpenStudy (anonymous):

a tug boat goes 120 miles upstream in 15 hours but downstream in 5 hours find the speed of the tugboas is ? mph and current speed is ?

hero (hero):

If you want to help @Chlorophyll, go ahead

OpenStudy (anonymous):

@Hero I'm not great at explanation. I'd rather to shadow you :)

OpenStudy (anonymous):

can someone please help me I have to ace this test

hero (hero):

I'm about to go to bed

OpenStudy (anonymous):

5 ( v + c) = 120 15( v - c) = 120 @bigmommatiad Can you solve it?

OpenStudy (anonymous):

no that is more confusing now would yo udivide 120 by 5 and then 120 by 15 to get the mph and current

OpenStudy (anonymous):

V = velocity, c = current

OpenStudy (anonymous):

5 ( v + c) = 120 downstream 15( v - c) = 120 upstream

OpenStudy (anonymous):

but how do I get what the v and c is that is where I am lost

OpenStudy (anonymous):

i thought that I would divide the 120 by 15 and the 120 by 5

OpenStudy (anonymous):

Yes -> v + c = 60 v - c = 8 Now add them together, ...?

OpenStudy (anonymous):

so the tugboat is 60 and the current is 8

OpenStudy (anonymous):

You don't know how to solve system of 2 unknowns ?

OpenStudy (anonymous):

omg I am going to flunk this test that I have been studying for all week

OpenStudy (anonymous):

to me no I do not think

OpenStudy (anonymous):

if I add them together it would be 68

OpenStudy (anonymous):

Add them VERTICALLY! 2V .....

OpenStudy (anonymous):

-> v = ....

OpenStudy (anonymous):

2v-2c=68

OpenStudy (anonymous):

+ C - C = ..?

OpenStudy (anonymous):

You have 1 candy, then you eat 1 candy, what left?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Assume C is candy! Now go back to the equation, after adding vertically: -> 2v .... = ?

OpenStudy (anonymous):

2v=68 then you would divide 68 b 2 to get v correct?

OpenStudy (anonymous):

Please do so, por favor :)

OpenStudy (anonymous):

so that would be v=34

OpenStudy (anonymous):

Yes, now plug v = 34 into v - c = 8 => c = ...?

OpenStudy (anonymous):

34-c=8=> c=42 dont you add the 34 to both sides or subtract it

OpenStudy (anonymous):

v - c = 8 v - 8 = c

OpenStudy (anonymous):

so that would be 34-8 which is 26 for c.....so the tugboat is 34 and the current is 26?

OpenStudy (anonymous):

Test your results by plugging back !

OpenStudy (anonymous):

omg so if I started at 68 then that means that the tugboat is 68 and the current is 26 please say I am

OpenStudy (anonymous):

sorry I am just freaking out on this test

OpenStudy (anonymous):

what does that mean

OpenStudy (anonymous):

Look what you wrote "2v=68 then you would divide 68 b 2 to get v correct? "

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

then " so that would be v=34 "

OpenStudy (anonymous):

so that would mean that 68 is tugboat and 34 would be current?

OpenStudy (anonymous):

V = velocity of tugboat , c = current

OpenStudy (anonymous):

v=tugboat and that would be 34 mph and current would be 0?

OpenStudy (anonymous):

no the current would be 26

OpenStudy (anonymous):

You also wrote " so that would be 34-8 which is 26 for c.....so the tugboat is 34 and the current is 26? "

OpenStudy (anonymous):

wow that so much are you going to be here longer cuz I am sure I am going to need more help.

OpenStudy (anonymous):

You definitely need to review this solution couple of dozens times to understand how to solve both ONE unknown and 2 unknowns. Good luck with your test :)

OpenStudy (anonymous):

do have a question if 120/15 is 8 how come the 120/5 is not 24 instead of 60

OpenStudy (anonymous):

@bigmommatiad Great job, I'm glad that you've caught my mistake =) So now you can practice solving it correctly: 5 ( v + c) = 120 downstream 15( v - c) = 120 upstream -> v + c = 24 -> v - c = 8